The average dividend yield of a random sample of 25 JSE listed companies this year was found to be 14.5%, with a sample standard deviation of 3.4%. Assume that dividend yields are normally distributed.
(a) Calculate, with 90% confidence, the actual mean dividend yield of all JSE listed companies this year. Interpret finding and show all steps.
(b) Calculate, with 95% confidence, the actual mean dividend yield of all JSE listed companies this year. Compare the interval with one calculated in 3.1.1 and show all the steps.

Answer by Boreal(15235)

Half-interval for 90%t-interval is t(0.95,df=24)*s/ssqrt(n)
=1.711*3.4/5=1.16%
(13.34%, 15.66%)
-
For half-interval of 95% more confidence will be a wider interval: t(0.975, df=24)*s/sqrt(n)
=2.064*3.4/5=1.40
(13.1%, 15.9%)